Solve for x.
3x^2 - 11x = - 6
(If there is more than one solution, separate them with commas.)

Answer by nerdybill(7384)

3x^2 - 11x = - 6
3x^2 - 11x + 6 = 0
factor the left:
(3x-2)(x-3) = 0
x = {2/3, 3}